OpenStudy (anonymous):

A light ray is traveling from medium A to medium B. As it enters medium B, it bends toward the normal. What can you say about the speed of light in the medium? The speed of light in medium A is more than that in medium B. The speed of light in medium A is less than that in medium B. The speed of light in medium A is the same as in medium B. The speed of light in medium B is zero.

OpenStudy (anonymous):

v = c / n and Snell's Law:\[n_{A}sin(\theta _{A})=n_{B}sin(\theta _{B})\]

OpenStudy (ookawaiioo):

If its bends towards the normal = it has a lower index of refraction(n) As I remember it, more bending, higher index of refraction(n)